Given the following information: Demand: Qd = 200 – 5P Supply: Qs = 5P

If a quantity tax of $2 per unit sold is imposed, Calculate:

(1)Demand and Supply equation after the tax.

(2)Buyer's price after tax

(3)Seller's price after tax

(4)Quantity after tax
Answer #1

Answer

(1)Demand and Supply equation after the tax.

converting the supply curve to inverse supply

Q=5P

5P=Q

P=0.2Q

adding the tax

P=0.2Q+2

converting it back

P=0.2Q+2

0.2Q=P-2

Q=5P-10

the supply equation after tax is:

Q=5P-10

and the demand equation will not change

Q=200-5P

(2)Buyer's price after tax

the equilibrium is at Qd=new Qs

200-5P=5P-10

10P=210

P=21

the price buyer pay is $21

(3)Seller's price after tax

Q=200-5*21=95 units

using the old supply equation

Q=5P

95=5P

P=19

the price producer receives is $19

(4)Quantity after tax

Q=200-5*21=95 units

the quantity is 95 units